# LeGrad: An Explainability Method for Vision Transformers via Feature Formation Sensitivity

### v1

@@ 10.2 Pertubation example
Figure 8 illustrates the perturbation-based benchmark of Section 4.1 in the main paper. Given the explainability map generated by the interpretatability method, for the negative (respectively positive) we progressively remove the most important (respectively the least important) part of the image. We then look at the decline in the model accuracy.
@@ 11 Personal and human subjects data
We acknowledge the use of datasets such as ImageNet and OpenImagesV7, which contain images sourced from the internet, potentially without the consent of the individuals depicted. We also recognize that the VL models used in this study were trained on the LAION-2B dataset, which may include sensitive content. We emphasize the importance of ethical considerations in the use of such data.

### v2

@@ 8.2 Pertubation example
Figure 8 illustrates the perturbation-based benchmark of Section 4.1 in the main paper. Given the explainability map generated by the explainability method, for the negative (respectively positive) we progressively remove the most important (respectively the least important) part of the image. We then look at the decline in the model accuracy.
@@ 8.3 Qualitative examples
8.3.1 Ablations of ReLU and Layer Aggregation
9 Gradient Distribution over Layers
Figures 10 & 10 extend the ”gradient distribution over Layers” analysis conducted Section 4.7 to more backbone size and more set of weights .
In Figures 12 & 13 & 14 & 15 & 16 & 17 & 18 provide the gradient distribution over layer independetly for each class in PascalVOC. We observe that for most models and classes, the layers contributing the most were typically located towards the end of the ViT. An interesting exception to this trend was observed for the ’person’ class, which exhibited a higher sensitivity to the middle layers across all model sizes and weight sets. We hypothesize that this is due to the high frequency of the ’person’ class in the training data, enabling the ViT to identify the object early in the layer sequence, thereby triggering the activation in the middle layers.
Furthermore, we note that the most activated layer varied significantly depending on the class and the model. This variability was observed even between two models of the same size, such as ViT-L(openai) and ViT-L(metaclip). This observation underscores the rationale behind the LeGrad method’s approach of utilizing multiple layers, hence alleviating the need to select a specific layer, as the optimal choice would differ from model to model.
Figure 9: Gradient distribution over layers on PascalVOC for different pretrained ViT-L/14.
Figure 10: Gradient distribution over layers on PascalVOC for different pretrained ViT-B/16.
Method CSDC PC DC D SD TS
GradCAM 0.75 0.67 0.68 0.91 0.7 0.48
Rollout 0.86 0.8 0.82 0.8 0.76 0.
Chefer LRP 0.91 0.92 0.89 0.9 0.74 0.95
LeGrad 0.90 0.83 0.92 1. 0.77 0.97
Table 10: Evaluation of ViT-B/16 on the FunnyBirds Co-12 setup
10 Sanity Check
The Co-12 recipe[27] is a set of 12 properties for evaluating the explanation quality of explainability method for machine learing models. These properties provide a comprehensive framework for assessing how well one can explain the decision-making process of a model. In [22] proposed a dataset, called FunnyBirds, to evaluate explainability methods for visual models.
As a sanity check for the proposed LeGrad method, we follow the authors guidelines and evaluate on the provided ViT-B/16111https://github.com/visinf/funnybirds-framework using LeGrad, showing improvement over gradient-based methods while being on par with LRP.
@@ 11 Personal and human subjects data
We acknowledge the use of datasets such as ImageNet and OpenImagesV7, which contain images sourced from the internet, potentially without the consent of the individuals depicted. We recognize that the VL models used in this study were trained on the LAION-2B dataset, which may include sensitive content. We emphasize the importance of ethical considerations in the use of such data.
Figure 12: Gradient Distributlion over Layers for different classes dataset for Laion2B-ViT-B/16.
Figure 13: Gradient Distributlion over Layers for different classes dataset for OpenAI-ViT-B/16.
Figure 14: Gradient Distributlion over Layers for different classes dataset for Laion400M-ViT-B/16.
Figure 15: Gradient Distributlion over Layers for different classes dataset for MetaCLIP-ViT-B/16.
Figure 16: Gradient Distributlion over Layers for different classes dataset for OpenAI-ViT-L/14.
Figure 17: Gradient Distributlion over Layers for different classes dataset for Laion400M-ViT-L/14.
Figure 18: Gradient Distributlion over Layers for different classes dataset for Laion2B-ViT-H/14.